OpenStudy (anonymous):

What is the value of the expression? 16 – 4 + (160 ÷ 20)4

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

I got 44

OpenStudy (anonymous):

(160/20)=8 so 16-4+(8)4 (8)4 = 32 16-4+32 = 44

Nnesha (nnesha):

again 4 is an exponent don't forget!

Nnesha (nnesha):

\[\huge\rm 16 – 4 + (160 ÷ 20)^4\] is this ur question right ?

OpenStudy (anonymous):

yes and I need help again

Nnesha (nnesha):

use ^ for exponents 2 to the 3rd power 2^3 like this

Nnesha (nnesha):

16 – 4 + (160 ÷ 20)^4 like this

Nnesha (nnesha):

alright start with the parentheses

OpenStudy (anonymous):

can u help me on the one I just posted please and ty

Nnesha (nnesha):

for this one answer isn't 44